# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

### v1

@@ E.2 Tiny Shakespeare Nano-GPT Teacher to Pico-GPT Student
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✗✗✓ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
Appendix F Vision Results
F.1 TinyImageNet
@@
F.1.1 ResNet50
Findings:
For the ResNet50 on TinyImageNet, we observe that the teacher seeds, Table 25, obtain a low train loss of 0.001 and a train accuracy of 0.99. This train performance coincides with a test accuracy of circa 0.60, resulting in a generalisation gap of circa 0.39.
For an alpha of 0.1, Table 29, we observe no significant knowledge transfer across all metrics except for Rank Disagreement with teacher seed 0. It has statistically significant transfer, but the increased similarity is extremely marginal, as observed with SIDDO and KD 0.1 having the same value to 3 significant figures, see Table 26. With this, we see a marginal prediction agreement of less than 0.5% for correct and incorrect predictions across teacher seeds, Figure 23(a). For alpha 0.5 and 0.9, we observe significant knowledge transfer for all bar Prediction Disagreement with alpha of 0.5 and 0.9 for teacher seed 2. However, this transfer is marginal, Tables 26, 27 and 28, and we observe a prediction agreement of less than 0.5% for correct and incorrect predictions across teacher seeds, Figure 17(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.001426 | 0.999800 | 2.070590 | 0.605300 |
| 1 | 0.001393 | 0.999800 | 2.051494 | 0.607900 |
| 2 | 0.001436 | 0.999800 | 2.051024 | 0.610600 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
@@ F.1.1 ResNet50
| KD 0.1 | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
F.1.2 ResNet50 with RandAugment
Findings:
For the ResNet50 on TinyImageNet with RandAugment, we observe that the teacher seeds, Table 25, obtain a high train loss and a train accuracy of circa 0.84. This train performance coincides with a test accuracy of circa 0.64, resulting in a generalisation gap of circa 0.2.
We observe significant knowledge transfer for all alpha values with a strong asymmetric transfer of knowledge favouring incorrect predictions as shown in Table 34 and Figure 26(a), respectively. However, it is important to note that despite significant and substantial knowledge transfer, we do not see any improvement in test accuracy over the control and random controls.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.672748 | 0.840410 | 1.620552 | 0.638800 |
| 1 | 0.678245 | 0.839200 | 1.629393 | 0.641800 |
| 2 | 0.667570 | 0.840750 | 1.624969 | 0.641100 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
@@ F.1.3 VGG19
For the VGG19 on the TinyImageNet, we observe a low train loss of circa and a train accuracy of 0.9998. As expected, given our results and discussion in the main body of the paper on the ResNet50, we see no significant transfer until an alpha of 0.9. With teacher seed 0 and 2 with an alpha of 0.9, we record significant transfer for Activation Distance and for teacher seed 0 on JS Divergence, as seen in Table 39. When we observe knowledge transfer with an alpha of 0.9, we observe a slight preference for positive agreement of test prediction; however, the results have a large SEM, and the amount of agreement is less than 0.5%, making the results less reliable and insignificant in either transfer direction.
@@ F.1.4 VGG19 with RandAugment
For the VGG19 on the TinyImageNet with RandAugment, we observe a high train loss of circa and a train accuracy of circa 0.93. As expected, given the results on the RandAugment ResNet50 that we present in the main body of the paper, we see substantial transfer across all alpha values; see Tables 41, 42, 43 and 44. This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 32(a).
@@ F.2 CIFAR10
All CIFAR10 architectures are trained with Adam optimiser with a learning rate of 0.001 and a batch size of 256 for 100 epochs. All data is normalised with a mean of 0.5 and a standard deviation of 0.5. The student vision architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. As aligned with all experiments we conduct, this is repeated for the three teachers trained on seeds 0-2.
Justification:
This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
Findings:
We find that the teacher models often significantly transfer knowledge to the student model, and this coincides with the teacher’s high loss on the training dataset. The ResNet has the lowest loss and no transfer, the VGG has a higher loss and some transfer, and the ViT has the highest loss and the most transfer. However, when knowledge is transferred, it often has a negative asymmetric payoff towards agreement between the teacher and the student on incorrect predictions.
F.2.1 ResNet18
Findings:
For the ResNet18 on CIFAR10, we observe that the teacher seeds, Table 45, obtain a very low train loss of and a train accuracy of 1. This train performance coincides with a high test accuracy of circa 0.86, resulting in a generalisation gap of circa 0.14. Table 49 shows no significant knowledge transfer across teacher seeds.
Due to the low train loss on the teacher seed, the teacher model is a nearly identical representation of the training labels, meaning there is low utility in the teacher model. As we observe, the controls of the models trained in the SIDDO condition is functionally different from the teacher, Tables 46, 47 and 48; despite having the same initialisation and only changing the data order, it is not a surprise that Knowledge Distillation in the setup does not add anything as the teacher is essentially the label, and thus creates a similar setup to the SIDDO condition.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000010 | 1.000000 | 0.869184 | 0.862100 |
| 1 | 0.000006 | 1.000000 | 0.833735 | 0.867200 |
| 2 | 0.000030 | 1.000000 | 0.739927 | 0.867000 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.174 | 0.175 | 0.172 | 0.174 | 0.244 | 0.538 | 0.843 |
| Rank Disagreement () | 0.659 | 0.659 | 0.656 | 0.655 | 0.795 | 0.802 | 0.807 |
| Prediction Disagreement () | 0.128 | 0.129 | 0.127 | 0.128 | 0.131 | 0.143 | 0.150 |
| JS Divergence () | 0.070 | 0.070 | 0.069 | 0.068 | 0.097 | 0.229 | 0.432 |
| Accuracy () | 0.861 | 0.862 | 0.862 | 0.862 | 0.865 | 0.856 | 0.854 |
| Loss () | 0.961 | 0.903 | 0.895 | 0.827 | 0.539 | 0.902 | 1.772 |
| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
| Activation Distance () | 0.167 | 0.164 | 0.165 | 0.165 | 0.240 | 0.533 | 0.841 |
| Rank Disagreement () | 0.653 | 0.649 | 0.650 | 0.650 | 0.796 | 0.803 | 0.807 |
| Prediction Disagreement () | 0.122 | 0.120 | 0.121 | 0.120 | 0.126 | 0.134 | 0.139 |
| JS Divergence () | 0.066 | 0.065 | 0.065 | 0.064 | 0.095 | 0.226 | 0.430 |
| Accuracy () | 0.865 | 0.867 | 0.866 | 0.867 | 0.866 | 0.860 | 0.859 |
| Loss () | 0.858 | 0.877 | 0.824 | 0.816 | 0.533 | 0.896 | 1.767 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.166 | 0.169 | 0.167 | 0.172 | 0.242 | 0.533 | 0.839 |
| Rank Disagreement () | 0.646 | 0.647 | 0.638 | 0.646 | 0.799 | 0.803 | 0.805 |
| Prediction Disagreement () | 0.122 | 0.124 | 0.124 | 0.127 | 0.132 | 0.140 | 0.142 |
| JS Divergence () | 0.065 | 0.066 | 0.064 | 0.067 | 0.096 | 0.226 | 0.429 |
| Accuracy () | 0.865 | 0.864 | 0.864 | 0.861 | 0.862 | 0.857 | 0.857 |
| Loss () | 0.892 | 0.887 | 0.803 | 0.798 | 0.549 | 0.900 | 1.769 |
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
F.2.2 VGG19
Findings:
For the VGG19 on CIFAR10, we observe that the teacher seeds, Table 50, obtain a low train loss of circa 0.01 and a train accuracy of approximately 0.996. This train performance coincides with a high test accuracy of circa 0.86, resulting in a generalisation gap of circa 0.14. Table 54 shows a significant knowledge transfer with regard to Rank Disagreement for all teacher seeds when alpha is at 0.9.
At alpha 0.9 for teacher seed 0 and 2, there is an increase in agreement between the student and teacher on incorrect predictions over the correct predictions, Figure 38(a), which corresponds with the knowledge transfer. This result coincides with teachers seed 0 and 2 having a higher train loss than teacher seed 1, indicating that the teacher train loss plays an important role in knowledge transfer. For teacher seed 1, Figure 38(a), there is no significant increase in correct or incorrect prediction agreement between the student model and the teacher due to the deviation in the SEM.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.011608 | 0.996760 | 0.858675 | 0.863900 |
| 1 | 0.009228 | 0.997080 | 0.798530 | 0.860800 |
| 2 | 0.012352 | 0.996420 | 0.801562 | 0.867100 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.206 | 0.199 | 0.203 | 0.197 | 0.264 | 0.541 | 0.842 |
| Rank Disagreement () | 0.701 | 0.705 | 0.658 | 0.640 | 0.811 | 0.819 | 0.819 |
| Prediction Disagreement () | 0.152 | 0.147 | 0.151 | 0.146 | 0.148 | 0.146 | 0.150 |
| JS Divergence () | 0.090 | 0.085 | 0.086 | 0.083 | 0.109 | 0.230 | 0.429 |
| Accuracy () | 0.864 | 0.869 | 0.867 | 0.869 | 0.870 | 0.871 | 0.868 |
| Loss () | 0.849 | 0.725 | 0.676 | 0.649 | 0.562 | 0.880 | 1.762 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.199 | 0.202 | 0.202 | 0.201 | 0.263 | 0.543 | 0.842 |
| Rank Disagreement () | 0.726 | 0.684 | 0.662 | 0.639 | 0.803 | 0.801 | 0.810 |
| Prediction Disagreement () | 0.147 | 0.150 | 0.150 | 0.149 | 0.148 | 0.149 | 0.153 |
| JS Divergence () | 0.086 | 0.087 | 0.086 | 0.085 | 0.107 | 0.230 | 0.428 |
| Accuracy () | 0.868 | 0.866 | 0.865 | 0.866 | 0.870 | 0.869 | 0.866 |
| Loss () | 0.799 | 0.735 | 0.680 | 0.666 | 0.562 | 0.887 | 1.762 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.196 | 0.199 | 0.196 | 0.193 | 0.258 | 0.541 | 0.844 |
| Rank Disagreement () | 0.672 | 0.649 | 0.633 | 0.602 | 0.809 | 0.817 | 0.816 |
| Prediction Disagreement () | 0.142 | 0.146 | 0.143 | 0.141 | 0.142 | 0.143 | 0.149 |
| JS Divergence () | 0.084 | 0.086 | 0.083 | 0.081 | 0.106 | 0.229 | 0.429 |
| Accuracy () | 0.870 | 0.864 | 0.868 | 0.867 | 0.871 | 0.871 | 0.867 |
| Loss () | 0.801 | 0.734 | 0.665 | 0.639 | 0.560 | 0.884 | 1.762 |
| KD 0.1 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✓✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ |
F.2.3 ViT
Findings:
For the ViT on CIFAR10, we observe that the teacher seeds, Table 55, obtain a high train loss of and a train accuracy of approximately 0.98. This train performance coincides with a test accuracy of circa 0.63, resulting in a generalisation gap of circa 0.35. Table 59 shows a significant knowledge transfer on all teacher seeds when alpha is 0.5 and 0.9. For teacher seed 0 and 1 using alpha at 0.9, where there is sizeable knowledge transfer, we observe an asymmetric knowledge transfer favouring negative transfer in Figure 41(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.043291 | 0.988260 | 1.864339 | 0.626900 |
| 1 | 0.056539 | 0.983160 | 1.772490 | 0.634200 |
| 2 | 0.046902 | 0.987100 | 1.714442 | 0.649600 |
| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
| Activation Distance () | 0.491 | 0.487 | 0.473 | 0.470 | 0.496 | 0.611 | 0.793 |
| Rank Disagreement () | 0.734 | 0.730 | 0.724 | 0.722 | 0.808 | 0.812 | 0.817 |
| Prediction Disagreement () | 0.385 | 0.383 | 0.374 | 0.373 | 0.383 | 0.380 | 0.386 |
| JS Divergence () | 0.201 | 0.198 | 0.189 | 0.186 | 0.206 | 0.277 | 0.411 |
| Accuracy () | 0.634 | 0.634 | 0.641 | 0.637 | 0.640 | 0.641 | 0.627 |
| Loss () | 1.773 | 1.695 | 1.52 | 1.451 | 1.258 | 1.351 | 1.943 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.485 | 0.477 | 0.461 | 0.455 | 0.489 | 0.609 | 0.791 |
| Rank Disagreement () | 0.733 | 0.728 | 0.717 | 0.714 | 0.806 | 0.808 | 0.816 |
| Prediction Disagreement () | 0.382 | 0.375 | 0.367 | 0.363 | 0.379 | 0.380 | 0.382 |
| JS Divergence () | 0.198 | 0.193 | 0.182 | 0.178 | 0.202 | 0.275 | 0.410 |
| Accuracy () | 0.637 | 0.643 | 0.644 | 0.648 | 0.643 | 0.636 | 0.630 |
| Loss () | 1.781 | 1.668 | 1.466 | 1.366 | 1.253 | 1.359 | 1.942 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.476 | 0.468 | 0.459 | 0.456 | 0.486 | 0.612 | 0.797 |
| Rank Disagreement () | 0.730 | 0.725 | 0.720 | 0.718 | 0.806 | 0.811 | 0.817 |
| Prediction Disagreement () | 0.372 | 0.366 | 0.363 | 0.360 | 0.371 | 0.374 | 0.375 |
| JS Divergence () | 0.195 | 0.189 | 0.183 | 0.180 | 0.201 | 0.277 | 0.413 |
| Accuracy () | 0.636 | 0.641 | 0.644 | 0.639 | 0.637 | 0.635 | 0.631 |
| Loss () | 1.788 | 1.673 | 1.498 | 1.458 | 1.282 | 1.361 | 1.942 |
| KD 0.1 | ✗✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
F.3 SVHN Dataset
Training Settings:
All SVHN architectures are trained with Adam optimiser with a learning rate of 0.001 and a batch size of 256 for 100 epochs. All data is normalised with a mean of 0.5 and a standard deviation of 0.5. The student vision architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. We repeated this, in line with our other experiments for the three teachers trained on seeds 0-2.
Justification:
This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
Findings:
The teacher models often significantly transfer knowledge to the student model. However, the knowledge transfer is often inconsistent, and when transferred, it often has an asymmetric negative payoff.
F.3.1 ResNet18
Findings:
For the ResNet on SVHN, we observe that the teacher seeds, Table 60, obtain a range of train loss values of 0.000646, 0.000061 and 0.004657 for teacher seeds 0, 1, and 2, respectively. The train accuracies are approximately 0.99. This train performance coincides with a test accuracy of circa 0.95, resulting in a generalisation gap of circa 0.04.
The teacher model with a higher training loss (seed 2) has significant knowledge transfer, see Table 64, for all functional similarity metrics across alpha values 0.1, 0.5 and 0.9, except for Prediction Disagreement when alpha was 0.1. In this case, we also observe a large asymmetric payoff in prediction agreement, significantly favouring incorrect predictions, Figure 44(a). Whereas teacher seed 0 has a train loss of 0.000061 and has no significant transfer with alpha values of 0.1 and 0.5. However, with an alpha of 0.9, it does have a significant transfer across metrics except for Prediction Disagreement, see Table 64. When alpha is 0.9, we observe an asymmetric payoff in prediction agreement, significantly favouring incorrect predictions. For teacher seed 0, which has a train loss of 0.000646, we observe significant knowledge transfer when alpha is 0.5 and 0.9, coinciding with an asymmetric payoff in prediction agreement, favouring incorrect predictions.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000646 | 0.999850 | 0.381410 | 0.951829 |
| 1 | 0.000061 | 0.999973 | 0.331054 | 0.952251 |
| 2 | 0.004657 | 0.998580 | 0.309702 | 0.947104 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| JS Divergence () | 0.025 | 0.025 | 0.023 | 0.022 | 0.053 | 0.201 | 0.431 |
| Accuracy () | 0.952 | 0.951 | 0.954 | 0.954 | 0.957 | 0.957 | 0.955 |
| Loss () | 0.385 | 0.344 | 0.310 | 0.293 | 0.236 | 0.692 | 1.698 |

### v2

@@ C.3 TinyShakespeare Nano-GPT Teacher to Pico-GPT Student
C.3.1 The Effect Of Temperature
This section provides the raw results for temperature 2 and 4 on the TinyShakespeare dataset.
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 0.000 | 0.197 0.000 | 0.183 0.000 | 0.181 0.000 | 0.213 0.000 | 0.305 0.001 | 0.617 0.000 |
| Rank Disagreement | 0.915 0.000 | 0.907 0.000 | 0.896 0.000 | 0.892 0.000 | 0.94 0.000 | 0.945 0.000 | 0.95 0.000 |
| Prediction Disagreement | 0.252 0.000 | 0.25 0.000 | 0.235 0.000 | 0.23 0.000 | 0.252 0.000 | 0.253 0.000 | 0.27 0.000 |
| JS Divergence | 0.056 0.000 | 0.053 0.000 | 0.047 0.000 | 0.047 0.000 | 0.072 0.000 | 0.152 0.000 | 0.403 0.000 |
| Accuracy | 0.571 0.000 | 0.572 0.000 | 0.572 0.000 | 0.569 0.000 | 0.571 0.000 | 0.571 0.000 | 0.562 0.000 |
| Loss | 1.473 0.002 | 1.513 0.003 | 1.571 0.002 | 1.622 0.002 | 1.493 0.001 | 1.736 0.001 | 2.732 0.001 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.201 0.000 | 0.195 0.000 | 0.181 0.000 | 0.179 0.000 | 0.209 0.000 | 0.298 0.000 | 0.609 0.000 |
| Rank Disagreement | 0.916 0.000 | 0.907 0.000 | 0.896 0.000 | 0.892 0.000 | 0.94 0.000 | 0.945 0.000 | 0.95 0.000 |
| Prediction Disagreement | 0.258 0.001 | 0.254 0.000 | 0.24 0.000 | 0.236 0.000 | 0.256 0.000 | 0.258 0.000 | 0.279 0.000 |
| JS Divergence | 0.055 0.000 | 0.052 0.000 | 0.047 0.000 | 0.046 0.000 | 0.071 0.000 | 0.15 0.000 | 0.401 0.000 |
| Accuracy | 0.571 0.000 | 0.571 0.000 | 0.572 0.000 | 0.569 0.000 | 0.572 0.000 | 0.571 0.000 | 0.56 0.000 |
| Loss | 1.474 0.002 | 1.512 0.003 | 1.569 0.002 | 1.613 0.003 | 1.489 0.001 | 1.732 0.001 | 2.739 0.001 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.201 0.000 | 0.195 0.000 | 0.181 0.000 | 0.18 0.000 | 0.21 0.000 | 0.301 0.000 | 0.615 0.000 |
| Rank Disagreement | 0.915 0.000 | 0.906 0.000 | 0.896 0.000 | 0.892 0.000 | 0.94 0.000 | 0.945 0.000 | 0.95 0.000 |
| Prediction Disagreement | 0.251 0.001 | 0.247 0.000 | 0.235 0.000 | 0.23 0.000 | 0.249 0.000 | 0.252 0.000 | 0.274 0.000 |
| JS Divergence | 0.055 0.000 | 0.052 0.000 | 0.046 0.000 | 0.046 0.000 | 0.071 0.000 | 0.15 0.000 | 0.403 0.000 |
| Accuracy | 0.571 0.000 | 0.571 0.000 | 0.571 0.000 | 0.569 0.000 | 0.572 0.000 | 0.571 0.000 | 0.56 0.000 |
| Loss | 1.474 0.002 | 1.513 0.001 | 1.576 0.001 | 1.619 0.003 | 1.489 0.001 | 1.732 0.001 | 2.739 0.001 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 0.000 | 0.199 0.000 | 0.189 0.000 | 0.193 0.000 | 0.206 0.000 | 0.262 0.000 | 0.568 0.001 |
| Rank Disagreement | 0.915 0.000 | 0.893 0.000 | 0.88 0.000 | 0.876 0.000 | 0.94 0.000 | 0.945 0.000 | 0.951 0.000 |
| Prediction Disagreement | 0.252 0.000 | 0.251 0.001 | 0.244 0.000 | 0.245 0.000 | 0.253 0.000 | 0.253 0.000 | 0.27 0.000 |
| JS Divergence | 0.056 0.000 | 0.054 0.000 | 0.05 0.000 | 0.051 0.000 | 0.067 0.000 | 0.127 0.000 | 0.362 0.000 |
| Accuracy | 0.571 0.000 | 0.57 0.000 | 0.568 0.000 | 0.562 0.000 | 0.572 0.000 | 0.571 0.000 | 0.562 0.000 |
| Loss | 1.473 0.002 | 1.528 0.002 | 1.592 0.002 | 1.663 0.002 | 1.491 0.002 | 1.68 0.000 | 2.544 0.002 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.201 0.000 | 0.196 0.000 | 0.188 0.000 | 0.191 0.000 | 0.203 0.000 | 0.256 0.000 | 0.562 0.000 |
| Rank Disagreement | 0.916 0.000 | 0.893 0.000 | 0.88 0.000 | 0.876 0.000 | 0.94 0.000 | 0.945 0.000 | 0.951 0.000 |
| Prediction Disagreement | 0.258 0.001 | 0.256 0.000 | 0.25 0.000 | 0.249 0.000 | 0.256 0.000 | 0.258 0.000 | 0.278 0.000 |
| JS Divergence | 0.055 0.000 | 0.052 0.000 | 0.049 0.000 | 0.05 0.000 | 0.066 0.000 | 0.126 0.000 | 0.361 0.000 |
| Accuracy | 0.571 0.000 | 0.57 0.000 | 0.568 0.000 | 0.563 0.000 | 0.571 0.000 | 0.571 0.000 | 0.561 0.000 |
| Loss | 1.474 0.002 | 1.528 0.002 | 1.59 0.002 | 1.653 0.003 | 1.489 0.001 | 1.677 0.001 | 2.55 0.002 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.201 0.000 | 0.197 0.000 | 0.189 0.000 | 0.192 0.000 | 0.204 0.000 | 0.259 0.000 | 0.567 0.000 |
| Rank Disagreement | 0.915 0.000 | 0.893 0.000 | 0.879 0.000 | 0.876 0.000 | 0.94 0.000 | 0.945 0.000 | 0.951 0.000 |
| Prediction Disagreement | 0.251 0.001 | 0.25 0.001 | 0.245 0.001 | 0.245 0.000 | 0.25 0.001 | 0.253 0.000 | 0.275 0.000 |
| JS Divergence | 0.055 0.000 | 0.053 0.000 | 0.049 0.000 | 0.05 0.000 | 0.066 0.000 | 0.127 0.000 | 0.363 0.000 |
| Accuracy | 0.571 0.000 | 0.57 0.000 | 0.568 0.000 | 0.562 0.000 | 0.571 0.000 | 0.571 0.000 | 0.561 0.000 |
| Loss | 1.474 0.002 | 1.53 0.001 | 1.594 0.002 | 1.658 0.002 | 1.489 0.001 | 1.677 0.001 | 2.55 0.002 |
Appendix D Vision Results
D.1 TinyImageNet
@@
D.1.1 ResNet50
Here we present the results of the ResNet50 on TinyImageNet in the Self Distillation regime for teacher seeds 1 and 2 - the corresponding analysis can be found in the main body in Section 4.1.
@@ D.1.1 ResNet50
D.1.2 ResNet50 with RandAugment
Here we present the results of the ResNet50 on TinyImageNet using RandAugment for student and teacher in the Self Distillation regime for teacher seeds 1 and 2 - the corresponding analysis can be found in the main body in Section 4.1.
@@ D.1.3 VGG19
For the VGG19 on the TinyImageNet, we observe a low train loss of circa and a train accuracy of 0.9998. As expected, given our results and discussion in the main body of the paper on the ResNet50, we see no statistically supported until an of 0.9. With teacher seed 0 and 2 with an of 0.9, we record statistically supported transfer for Activation Distance and for teacher seed 0 on JS Divergence, as seen in Table 55. When we observe knowledge transfer with an of 0.9, we observe a slight preference for positive agreement of test prediction; however, the results have a large SEM, and the amount of agreement is less than 0.5%, making the results less reliable and insignificant in either transfer direction.
@@ D.1.4 VGG19 with RandAugment
For the VGG19 on the TinyImageNet with RandAugment, we observe a high train loss of circa and a train accuracy of circa 0.93. As expected, given the results on the RandAugment ResNet50 that we present in the main body of the paper, we see statistically supported transfer across all values; see Tables 57, 58, 59 and 60. This substantial and statistically supported transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 19.
@@ D.2 SVHN Dataset
All SVHN architectures are trained with Adam with a learning rate of 0.001 and a batch size of 256 for 100 epochs. All data is normalised with a mean of 0.5 and a standard deviation of 0.5. The student vision architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. We repeated this, in line with our other experiments for the three teachers trained on seeds 0-2.
D.2.1 ResNet18